# A Survey of Event Causality Identification: Principles, Taxonomy, Challenges, and Assessment

### v1

@@ Abstract
Event Causality Identification (ECI) has become a crucial task in Natural Language Processing (NLP), aimed at automatically extracting causalities from textual data. In this survey, we systematically address the foundational principles, technical frameworks, and challenges of ECI, offering a comprehensive taxonomy to categorize and clarify current research methodologies, as well as a quantitative assessment of existing models. We first establish a conceptual framework for ECI, outlining key definitions, problem formulations, and evaluation standards. Our taxonomy classifies ECI methods according to the two primary tasks of sentence-level (SECI) and document-level (DECI) event causality identification. For SECI, we examine feature pattern-based matching, deep semantic encoding, causal knowledge pre-training and prompt-based fine-tuning, and external knowledge enhancement methods. For DECI, we highlight approaches focused on event graph reasoning and prompt-based techniques to address the complexity of cross-sentence causal inference. Additionally, we analyze the strengths, limitations, and open challenges of each approach. We further conduct an extensive quantitative evaluation of various ECI methods on two benchmark datasets. Finally, we explore future research directions, highlighting promising pathways to overcome current limitations and broaden ECI applications.
@@ 1 Introduction
The remainder of this paper is organized as follows: Section 2 presents the relevant concepts of ECI, reviews commonly used datasets, and evaluation metrics. Section 3 introduces key technologies for ECI. Section 4 provides a comprehensive overview of the classification framework for ECI, summarizes core techniques of various models, and analyzes their strengths and weaknesses. Section 5 presents a quantitative evaluation of the performance of existing classical methods on two datasets. Section 6 outlines future research directions for ECI.
@@ 3.3.3 Graph Embedding
4 Taxonomy
This section distinguishes SECI from DECI, presenting relevant methods and models. SECI and DECI are reviewed separately due to differences in task complexity, model demands, and challenges, with DECI being the more difficult. Methods are first categorized by their characteristics, followed by a discussion of their core techniques, classic models, strengths, and open challenges.
4.1 Sentence-Level Event Causality Identification
@@
4.1.1 Feature Pattern-Based Matching
@@
Open Challenges. Pattern matching-based methods are simple, efficient, and interpretable, with low computational demands, making them popular in early research. However, they lack the ability to capture deep semantic information and rely heavily on external parsers and predefined lexical or syntactic patterns. This dependence limits their scalability, flexibility, and generalizability, causing performance to degrade significantly with complex semantics, unfamiliar structures, or noisy text.
4.1.2 Deep Semantic Encoding
@@
4.1.2.1 Simple Encoding Strategy
@@
$\mathbf{p}=\mathrm{MLP}(\mathbf{ep}_{ij})=\mathrm{softmax}(\mathbf{W\cdot ep}_{ij}+\mathbf{b}),$ (9)
$\mathbf{W}$ and $\mathbf{b}$ are learnable parameter matrices, and $\mathbf{p}\in\mathbb{R}^{2}$ (ignoring direction) or $\mathbf{p}\in\mathbb{R}^{3}$ (considering direction). Each dimension of $\mathbf{p}$ corresponds to the probability of the causality or non-causality.
@@
4.1.2.2 Enhanced Encoding Strategy
@@
Summary. Simple encoding strategies are easy to implement and train, offering high computational efficiency but are only suited for clear, low-noise texts within a single domain. Enhanced encoding strategies provide a more comprehensive modeling of event causality, improving accuracy and robustness in ECI, especially in complex contexts and long-distance dependencies. They are effective in handling complex semantic structures and exhibit stronger cross-domain generalizability. However, deep semantic encoding-based methods require a substantial amount of labeled data. Without sufficient supervised learning, these models struggle to learn high-quality encoding and classification capabilities, which impacts inference accuracy.
4.1.3 Causal Knowledge Pretraining and Prompt-Based Fine-Tuning
@@
4.1.3.1 Causal Knowledge Pre-training
@@
4.1.3.2 Prompt-Based Fine-Tuning
@@ 4.1.3.3 Open Challenges
4.1.4 External Knowledge Enhancement
@@ 4.1.4.4 Open Challenges
4.2 Document-Level Event Causality Identification
@@
4.2.1 Event Graph Reasoning-Based Methods
@@
ERGO [112] Event Pairs Causalities
@@
4.2.2 Prompt-based methods
@@ 4.3 Performance Assessment
TABLE V: Evaluation includes five evaluation aspects and all the results are relative. "I" assesses the ability of a method to identify implicit causalities($\star$: Strong ability $\sqrt{}$: Moderate ability ×: Weak ability). "A" indicates the level of annotated data required for effective performance (×: High requirement $\sqrt{}$: Moderate requirement $\star$: Low requirement). "G" measures the method’s ability to generalize across different topics or domains ($\star$: Strong generalization capability $\sqrt{}$: Moderate capability ×: Weak capability). "E" indicates the degree to which a method relies on external knowledge (↑: High dependency →: Moderate dependency -: No dependency). "F" assesses the extent to which a method incorrectly reports false positives (×: Severe false $\sqrt{}$: Minimal false positives $\star$: Little to no false positives).
@@ 5 Experimental Evaluation
5.1 Experimental Setup
5.1.1 Dataset Processing
The details of the two datasets used in this experiment, Causal-TimeBank (CTB) and Event StoryLine v0.9 (ESL), are provided in Section 2.2. For consistency with prior studies, we adopted the common data splits: in the ESL dataset, the last two topics were designated as development data, while the remaining 20 topics were used for 5-fold cross-validation. For the CTB dataset, we conducted 10-fold cross-validation.
5.1.2 Evaluation Protocol
We evaluated the models using Precision (P), Recall (R), and F1-score (F1) as metrics, distinguishing between intra-sentence causality, inter-sentence causality, and overall performance. Given that most existing methods evaluate only the presence of causalities without considering directional accuracy, we focused our comparison similarly on causality presence alone. For the CTB dataset, where only 18 inter-sentence causal pairs exist, we limited evaluation to intra-sentence causality. For the ESL dataset, we assessed performance on intra-sentence causality, inter-sentence causality, and overall identification.
5.1.3 Baseline Models
We included several state-of-the-art models from existing literature, along with standard baselines like LSTM and BERT (used as encoding and classification frameworks). Additionally, we incorporated multiple LLMs as baselines, including GPT-3 (text-davinci), GPT-3.5-turbo, GPT-$4$https://openai.com/chatgpt/, and LLaMA2-7B  [^5] http://huggingface.co/meta-llama/Llmma-2-7b. For the LLMs, we used a standardized prompt:
@@ 5.1.3 Baseline Models
5.1.4 Implementation Details
To ensure fair comparisons, if baseline models had reported results on CTB and ESL with consistent data splits, we directly cited those results. For baselines where the original paper used different datasets or splits but provided open-source code, we re-evaluated the model by reproducing it according to our specified data splits. Reproduction strictly followed hyperparameters reported in the original paper or code to maintain result comparability. Models without open-source code or using entirely different datasets were excluded. Additionally, for models with multiple variants, we reported the variant with the highest F1-score.
5.2 Results and Analysis
@@
5.2.1 Baseline Performance on SECI Task
Feature pattern matching-based methods struggle with balancing precision and recall, resulting in limited overall performance. Shallow feature-based methods, such as DD and VR-C, show high precision but low recall, while simple statistical methods like PPMI underperform in both.
@@ 5.2.1 Baseline Performance on SECI Task
5.2.2 Baseline Performance on DECI Task
Feature pattern recognition methods, such as OP, LIP, and LR+, deliver relatively high recall but with many false positives. Simple encoding-based methods show low precision and recall due to difficulty capturing context and causal links, especially between distant events.

### v2

# A Survey of Event Causality Identification: Taxonomy, Challenges, Assessment, and Prospects

@@ Abstract
Event Causality Identification (ECI) has become a crucial task in Natural Language Processing (NLP), aimed at automatically identifying causalities between events in texts. In this survey, we systematically address the foundational principles and technical frameworks, offering a comprehensive taxonomy to categorize and clarify current research methodologies. We also discuss their corresponding challenges and provide a quantitative assessment. We first establish a conceptual framework for ECI, outlining key definitions, problem formulations, and evaluation standards. Our taxonomy classifies ECI methods according to the two primary tasks of sentence-level (SECI) and document-level (DECI) event causality identification. For SECI, we examine feature pattern-based matching, deep semantic encoding, causal knowledge pre-training and prompt-based fine-tuning, and external knowledge enhancement methods. For DECI, we highlight approaches focused on event graph reasoning and prompt-based techniques. Additionally, we analyze the strengths, limitations, and open challenges of each approach. We further conduct an extensive quantitative evaluation of various ECI methods on two benchmark datasets. Finally, we explore future research directions, highlighting promising pathways to overcome current limitations and broaden ECI applications.
@@ 1 Introduction
The remainder of this paper is organized as follows: Section 2 presents the relevant concepts of ECI, reviews commonly used datasets, and evaluation metrics. Section 3 introduces key technologies for ECI. Section 4 and 5 provide a comprehensive overview of the classification framework for ECI, summarizes core techniques of various models, and analyzes their strengths and weaknesses. Section 6 presents a quantitative evaluation of the performance of existing classical methods on two datasets. Section 7 outlines future research directions for ECI.
@@ 3.3.3 Graph Embedding
4 Sentence-Level Event Causality Identification
@@
4.1 Feature Pattern-Based Matching
@@ 4.1.4 Feature classification-based methods
4.1.5 Open Challenges
Pattern matching-based methods are simple, efficient, and interpretable, with low computational demands, making them popular in early research. However, they lack the ability to capture deep semantic information and rely heavily on external parsers and predefined lexical or syntactic patterns. This dependence limits their scalability, flexibility, and generalizability, causing performance to degrade significantly with complex semantics, unfamiliar structures, or noisy text.
4.2 Deep Semantic Encoding
@@
4.2.1 Simple Encoding Strategy
@@
$\mathbf{p}=\mathrm{softmax}(\mathrm{MLP}(\mathbf{ep}_{ij}))=\mathrm{softmax}(f\mathbf{(W\cdot ep}_{ij}+\mathbf{b})),$ (9)
$\mathbf{W}$ and $\mathbf{b}$ are learnable parameter matrices, $f$ is the activation function, and $\mathbf{p}\in\mathbb{R}^{2}$ (ignoring direction) or $\mathbf{p}\in\mathbb{R}^{3}$ (considering direction). Each dimension of $\mathbf{p}$ corresponds to the probability of the causality or non-causality.
@@
4.2.2 Enhanced Encoding Strategy
@@
4.2.3 Open Challenges
Simple encoding strategies are easy to implement and train, offering high computational efficiency but are only suited for clear, low-noise texts within a single domain. Enhanced encoding strategies provide a more comprehensive modeling of event causality, improving accuracy and robustness in ECI, especially in complex contexts and long-distance dependencies. They are effective in handling complex semantic structures and exhibit stronger cross-domain generalizability. However, deep semantic encoding-based methods require a substantial amount of labeled data. Without sufficient supervised learning, these models struggle to learn high-quality encoding and classification capabilities, which impacts inference accuracy.
4.3 Causal Knowledge Pretraining and Prompt-Based Fine-Tuning
@@
4.3.1 Causal Knowledge Pre-training
@@
4.3.2 Prompt-Based Fine-Tuning
@@ 4.3.3 Open Challenges
4.4 External Knowledge Enhancement
@@ 4.4.4 Open Challenges
5 Document-Level Event Causality Identification
@@
5.1 Event Graph Reasoning-Based Methods
@@
ERGO [112] Event pairs Shared events
@@
5.2 Prompt-based Fine-Tuning
@@ 6.1 Performance Assessment
TABLE V: Assessment on five aspects, all the results are relative. "I" assesses the ability of a method to identify implicit causalities($\star$: Strong ability $\sqrt{}$: Moderate ability ×: Weak ability). "A" indicates the level of annotated data required for effective performance (×: High requirement $\sqrt{}$: Moderate requirement $\star$: Low requirement). "G" measures the method’s ability to generalize across different topics or domains ($\star$: Strong generalization capability $\sqrt{}$: Moderate capability ×: Weak capability). "E" indicates the degree to which a method relies on external knowledge (↑: High dependency →: Moderate dependency -: No dependency). "F" assesses the extent to which a method incorrectly reports false positives (×: Severe false $\sqrt{}$: Minimal false positives $\star$: Little to no false positives).
@@ 6.2 Experimental Evaluation
6.2.1 Experimental Setup
Dataset Processing. The details of the two datasets used in this experiment, Causal-TimeBank (CTB) and Event StoryLine v0.9 (ESL), are provided in Section 2.2. For consistency with prior studies, we adopted the common data splits: in the ESL dataset, the last two topics were designated as development data, while the remaining 20 topics were used for 5-fold cross-validation. For the CTB dataset, we conducted 10-fold cross-validation. Evaluation Protocol. We evaluated the models using Precision (P), Recall (R), and F1-score (F1) as metrics, distinguishing between intra-sentence causality, inter-sentence causality, and overall performance. Given that most existing methods evaluate only the presence of causalities without considering directional accuracy, we focused our comparison similarly on causality presence alone. For the CTB dataset, where only 18 inter-sentence causal pairs exist, we limited evaluation to intra-sentence causality. For the ESL dataset, we assessed performance on intra-sentence causality, inter-sentence causality, and overall identification.
Baselines. We included several state-of-the-art models from existing literature, along with standard baselines like LSTM and BERT (used as encoding and classification frameworks). Additionally, we incorporated multiple LLMs as baselines, including GPT-3 (text-davinci), GPT-3.5-turbo, GPT-$4$https://openai.com/chatgpt/, and LLaMA2-7B  [^5] http://huggingface.co/meta-llama/Llmma-2-7b. For the LLMs, we used a standardized prompt:
@@ 6.2.1 Experimental Setup
Implementation Details. To ensure fair comparisons, if baseline models had reported results on CTB and ESL with consistent data splits, we directly cited those results. For baselines where the original paper used different datasets or splits but provided open-source code, we re-evaluated the model by reproducing it according to our specified data splits. Reproduction strictly followed hyperparameters reported in the original paper or code to maintain result comparability. Models without open-source code or using entirely different datasets were excluded. Additionally, for models with multiple variants, we reported the variant with the highest F1-score.
6.2.2 Results and Analysis
@@
Baseline Performance on SECI Task. Feature pattern matching-based methods struggle with balancing precision and recall, resulting in limited overall performance. Shallow feature-based methods, such as DD and VR-C, show high precision but low recall, while simple statistical methods like PPMI underperform in both.
@@
Baseline Performance on DECI Task. Feature pattern recognition methods, such as OP, LIP, and LR+, deliver relatively high recall but with many false positives. Simple encoding-based methods show low precision and recall due to difficulty capturing context and causal links, especially between distant events.